### v1

@@ 4.1 An Artificial Game and Equilibrium Convergence Analysis
For $(\gamma,t,\lambda,\mu)\in\mathbb{R}^{n_{0}}\times[0,1]\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, let $f_{1}(\gamma,t,\lambda,\nu)=(f_{1}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $f_{1}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a)=\rho(t)\eta^{0i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}}a)$, $f_{2}(\gamma,t,\lambda,\nu)=(f_{2}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $f_{2}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a)=-\lambda^{i}(w^{i}_{I^{j}_{i}}a)$. Furthermore, define $p_{1}(\gamma,t,\lambda,\nu)=(p_{1}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $p_{1}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a)=g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i})+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}$, $p_{2}(\gamma,t,\lambda,\nu)=(p_{2}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a):i\in N,(j,a)\notin D_{i})$ with $p_{2}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a)=g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i})-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)$, $p_{3}(\gamma,t,\lambda,\nu)=(p_{3}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $p_{3}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}a)=(\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\rho(t)\eta^{0i}(w^{i}_{I^{j}_{i}}a))\lambda^{i}(w^{i}_{I^{j}_{i}}a)$, $p_{4}(\gamma,t,\lambda,\nu)=(p_{4}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}}):i\in N,j\in M_{i})$ with $p_{4}(\gamma,t,\lambda,\nu;w^{i}_{I^{j}_{i}})=\sum_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}})$. Then there exist constants $\tau_{1}>0,\kappa_{1}>0$ and $\kappa^{\prime}_{1}>0$ such that
$dist((\gamma,t,\lambda,\nu),\mathscr{C}_{E})$ $=\operatorname*{min}_{(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma,t,\lambda,\nu)-(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\|$
$\leq\tau_{1}(1+\|(\gamma,t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}(\|[f_{1}(\gamma,t,\lambda,\nu)]_{+}\|+\|[f_{2}(\gamma,t,\lambda,\nu)]_{+}\|$
$+\|[p_{1}(\gamma,t,\lambda,\nu)]_{+}\|+\|p_{2}(\gamma,t,\lambda,\nu)\|+\|p_{3}(\gamma,t,\lambda,\nu)\|+\|p_{4}(\gamma,t,\lambda,\nu)\|)^{\kappa_{1}}$
@@
$\begin{array}[]{l}g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=c(t)g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t)),\;i\in N,(j,a)\in D_{i},\\ g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=c(t)g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t)),\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\gamma^{i}(t;w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\rho(t)\eta^{0i}(w^{i}_{I^{j}_{i}}a))\lambda^{i}(w^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a),\\ \hskip 105.2751pt\rho(t)\eta^{0i}(w^{i}_{I^{j}_{i}}a)<\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i}.\end{array}$ (11)
@@
$\begin{array}[]{l}\operatorname*{min}\limits_{(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma(t),t,\lambda,\nu)-(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\|\leq\tau_{1}(1+\|(\gamma(t),t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}c(t)^{\kappa_{1}}\\ \hskip 99.58464pt(\sum\limits_{i\in N}\sum\limits_{j\in M_{i},a\in A(I^{j}_{i})}\|g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))\|+\sum\limits_{i\in N}\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a))^{\kappa_{1}}.\end{array}$
@@
$\begin{array}[]{l}\tau_{L}=\tau_{1}\max\limits_{(\gamma(t),t,\lambda,\nu)\in\mathscr{C}^{R}_{L}}(1+\|(\gamma(t),t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}\\ \hskip 99.58464pt(\sum\limits_{i\in N}\sum\limits_{j\in M_{i},a\in A(I^{j}_{i})}\|g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))\|+\sum\limits_{i\in N}\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)))^{\kappa_{1}}.\end{array}$
The compactness of $\mathscr{C}^{R}_{L}$ demonstrated in Appendix A confirms that $\tau_{L}$ is finite. Thus $\min_{(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma(t),t,\lambda,\nu)-(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\|\leq\tau_{L}c(t)^{\kappa_{L}}$ for any $(\gamma(t),t,\lambda,\nu)\in\mathscr{C}^{R}_{L}$. The proof is completed. ∎
@@
Let $\{\gamma^{*}(t_{k})\}_{k=1}^{\infty}$ be a sequence of Nash equilibria defined by the system (10) with $t=t_{k}\in(0,1]$ and $\lim_{k\to\infty}t_{k}=0$. Then every limit point of the totally mixed strategy sequence $\{\sigma(\gamma^{*}(t_{k}))\}_{k=1}^{\infty}$ yields a normal-form perfect equilibrium.
@@
For each $\gamma^{*}(t_{k})$, let $(\lambda^{*}(t_{k}),\nu^{*}(t_{k}))$ be the associated pair satisfying the system (10). Due to the compactness of $\mathscr{C}^{R}_{L}$, the sequence $\{(\gamma^{*}(t_{k}),t_{k},\lambda^{*}(t_{k}),\nu^{*}(t_{k}))\}_{k=1}^{\infty}$ has a convergent subsequence, which we denote in the same notation. As to each $t_{k}$, assume that $(\tilde{\gamma}(\tilde{t}_{k}),\tilde{t}_{k},\tilde{\lambda}(\tilde{t}_{k}),\tilde{\nu}(\tilde{t}_{k}))\in\arg\min_{(\tilde{\gamma}(\tilde{t}),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma^{*}(t_{k}),t_{k},\lambda^{*}(t_{k}),\nu^{*}(t_{k}))-(\tilde{\gamma}(\tilde{t}),\tilde{t},\tilde{\lambda},\tilde{\nu})\|$. Lemma 5 implies that there exist constants $\tau_{L}>0$ and $\kappa_{L}>0$ such that $|\tilde{t}_{k}-t_{k}|\leq\tau_{L}c(t_{k})^{\kappa_{L}}$. Given that $c(t_{k})=\exp(1-1/\rho(t_{k}))$, a sufficiently large constant $K_{0}$ ensures $\tau_{L}c(t_{k})^{\kappa_{L}}<t_{k}$ when $k>K_{0}$. It follows that $\tilde{t}_{k}>0$ for all $k>K_{0}$. Consequently, we deduce that $\lim_{k\to\infty}\sigma(\gamma^{*}(t_{k}))=\lim_{k\to\infty}\sigma(\tilde{\gamma}(\tilde{t}_{k}))$, indicating a normal-form perfect equilibrium. This completes the proof. ∎
@@ 4.2 A Smooth Path to a Normal-Form Perfect Equilibrium
In this subsection, we prove the existence of a smooth path along which the points satisfy the system (10). This path starts from a realization plan and ultimately converge into a normal-form perfect equilibrium.
@@
At $t=2$, the system (10) has a unique solution, given by $(\gamma^{*}(2),\lambda^{*}(2),\nu^{*}(2))$, with the components satisfying $\gamma^{*i}(2;w^{i}_{I^{j}_{i}}a)=\gamma^{0i}(w^{i}_{I^{j}_{i}}a)$, $\lambda^{i}(2;w^{*i}_{I^{j}_{i}}a)=1$ and $\nu^{*i}_{I^{j}_{i}}(2)=1$.
@@
$\begin{array}[]{l}\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(2;w^{i}_{I^{j}_{i}}a)-\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(2;w^{i}_{I^{j}_{i}}a)\lambda^{i}(w^{i}_{I^{j}_{i}}a)=\gamma^{0i}(w^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ 0<\gamma^{i}(2;w^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (12)
Suppose that $(\gamma^{*}(2),\lambda^{*}(2),\nu^{*}(2))$ is a solution to the system (12). We proceed by substituting the first group of equations into the third group, and summing over $a\in A(I^{j}_{i})$ for each $i\in N,j\in M_{i}$, yielding $\nu^{*i}_{I^{j}_{i}}(2)=1$. Then the first group of equations implies that $\lambda^{*i}(2;w^{i}_{I^{j}_{i}}a)=\nu^{*i}_{I^{j}_{i}}(2)=1$ for $i\in N,j\in M_{i},a\in A(I^{j}_{i})$. Substituting this into the third group of equations, we can conclude that $\gamma^{*i}(2;w^{i}_{I^{j}_{i}}a)=\gamma^{0i}(w^{i}_{I^{j}_{i}}a)$ for $i\in N,j\in M_{i},a\in A(I^{j}_{i})$. The proof is complete. ∎
As shown in Lemma 6, the system (10) admits a unique solution at $t=2$. We then identify a connected component containing this solution that extends to intersect the $t=0$ level. To facilitate this, it is crucial to introduce Browder’s fixed-point theorem [42].
@@
Let $\Omega_{L}=\cup_{t\in(0,2]}\Omega(t)$. For $(\hat{\gamma},t)\in\Omega_{L}\times(0,2]$, define $\varphi(\hat{\gamma},t)=(\gamma^{i}(t):i\in N)$, where $\gamma^{i}(t)$ denotes the unique solution to the strictly convex optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-c(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}))\\ &+c(t)\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a))\\ &+\theta(t)\sum\limits_{(j,a)\notin D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;w^{i}_{I^{j}_{i}}a))-\frac{1}{2}\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\hat{\gamma}^{i}(w^{i}_{I^{j}_{i}}a))^{2}\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (13)
@@
For $(\hat{\gamma},0)\in\Omega_{L}\times\{0\}$, let $\varphi(\hat{\gamma},0)=(\gamma^{i}(0):i\in N)$, where $\gamma^{i}(0)$ denotes the unique solution to the strictly convex optimization problem,
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(0)}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(0;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i})-\frac{1}{2}\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\gamma^{i}(0;w^{i}_{I^{j}_{i}}a)-\hat{\gamma}^{i}(w^{i}_{I^{j}_{i}}a))^{2}\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(0;w^{i}_{I^{j}_{i}}a)-\gamma^{i}(0;w^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (14)
@@
Based on Theorem 2.2.2 in [43], it follows that $\varphi(\gamma,t)$ is a continuous function that maps from $\Omega_{L}\times[0,2]$ to $\Omega_{L}$. Let =$\mathscr{F}=\{(\gamma(t),t)\in\Omega_{L}\times[0,2]|\varphi({\gamma}(t),t)={\gamma}(t)\}$. Theorem 4 ensures the existence of a connected component in $\mathscr{F}$ that intersects both $\mathbb{R}^{n_{0}}\times\{2\}$ and $\mathbb{R}^{n_{0}}\times\{0\}$. We denote this connected component as $\mathscr{F}^{c}$, and specially refer to the portion meeting $t>0$ as ~$\widetilde{\mathscr{F}}^{c}$.

### v3

@@ 4.1 An Artificial Game and Equilibrium Convergence Analysis
For $(\gamma,t,\lambda,\mu)\in\mathbb{R}^{n_{0}}\times[0,1]\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, let $f_{1}(\gamma,t,\lambda,\nu)=(f_{1}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $f_{1}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a)=\rho(t)\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)$, $f_{2}(\gamma,t,\lambda,\nu)=(f_{2}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $f_{2}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a)=-\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)$. Furthermore, define $p_{1}(\gamma,t,\lambda,\nu)=(p_{1}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $p_{1}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a)=g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i})+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}$, $p_{2}(\gamma,t,\lambda,\nu)=(p_{2}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a):i\in N,(j,a)\notin D_{i})$ with $p_{2}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a)=g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i})-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)$, $p_{3}(\gamma,t,\lambda,\nu)=(p_{3}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a):i\in N,(j,a)\in D_{i})$ with $p_{3}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}a)=(\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\rho(t)\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)$, $p_{4}(\gamma,t,\lambda,\nu)=(p_{4}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}}):i\in N,j\in M_{i})$ with $p_{4}(\gamma,t,\lambda,\nu;\varpi^{i}_{I^{j}_{i}})=\sum_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}})$. Then there exist constants $\tau_{1}>0,\kappa_{1}>0$ and $\kappa^{\prime}_{1}>0$ such that
$\begin{array}[]{l}dist((\gamma,t,\lambda,\nu),\mathscr{C}_{E})=\operatorname*{min}_{(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma,t,\lambda,\nu)-(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\|\\ \hskip 42.67912pt\leq\tau_{1}(1+\|(\gamma,t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}(\|[f_{1}(\gamma,t,\lambda,\nu)]_{+}\|+\|[f_{2}(\gamma,t,\lambda,\nu)]_{+}\|\\ \hskip 56.9055pt+\|p_{1}(\gamma,t,\lambda,\nu)\|+\|p_{2}(\gamma,t,\lambda,\nu)\|+\|p_{3}(\gamma,t,\lambda,\nu)\|+\|p_{4}(\gamma,t,\lambda,\nu)\|)^{\kappa_{1}}\end{array}$
@@
$\begin{array}[]{l}g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=c(t)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t)),\;i\in N,(j,a)\in D_{i},\\ g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=c(t)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t)),\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\rho(t)\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a),\\ \hskip 48.36958pt\rho(t)\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a)<\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;0<\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i}.\end{array}$ (11)
@@
$\begin{array}[]{l}\operatorname*{min}\limits_{(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\in\mathscr{C}_{E}}\|(\gamma(t),t,\lambda,\nu)-(\tilde{\gamma}(t),\tilde{t},\tilde{\lambda},\tilde{\nu})\|\leq\tau_{1}(1+\|(\gamma(t),t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}c(t)^{\kappa_{1}}\\ \hskip 71.13188pt(\sum\limits_{i\in N}\sum\limits_{j\in M_{i},a\in A(I^{j}_{i})}\|g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))\|+\sum\limits_{i\in N}\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a))^{\kappa_{1}}.\end{array}$
@@
$\begin{array}[]{l}\tau_{L}=\tau_{1}\max\limits_{(\gamma(t),t,\lambda,\nu)\in\mathscr{C}^{R}_{L}}(1+\|(\gamma(t),t,\lambda,\nu)\|)^{\kappa_{1}^{\prime}}\\ \hskip 71.13188pt(\sum\limits_{i\in N}\sum\limits_{j\in M_{i},a\in A(I^{j}_{i})}\|g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))\|+\sum\limits_{i\in N}\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)))^{\kappa_{1}}.\end{array}$
The compactness of $\mathscr{C}^{R}_{L}$ demonstrated in Appendix A confirms that $\tau_{L}$ is finite. Thus
for any $(\gamma(t),t,\lambda,\nu)\in\mathscr{C}^{R}_{L}$. The proof is completed. ∎
@@
Let $\{\gamma^{*}(t_{k})\}_{k=1}^{\infty}$ be a sequence of Nash equilibria of $\Gamma_{s}^{l}(t)$ defined by the system (10) with $t=t_{k}\in(0,1]$ and $\lim_{k\to\infty}t_{k}=0$. Then every limit point of the totally mixed strategy sequence $\{\sigma(\gamma^{*}(t_{k}))\}_{k=1}^{\infty}$ yields a normal-form perfect equilibrium.
@@
For each $\gamma^{*}(t_{k})$, let $(\lambda^{*}(t_{k}),\nu^{*}(t_{k}))$ be the associated pair satisfying the system (10). Due to the compactness of $\mathscr{C}^{R}_{L}$, the sequence $\{(\gamma^{*}(t_{k}),t_{k},\lambda^{*}(t_{k}),\nu^{*}(t_{k}))\}_{k=1}^{\infty}$ has a convergent subsequence, which we denote in the same notation. As to each $t_{k}$, assume that
[…]
Lemma 5 implies that there exist constants $\tau_{L}>0$ and $\kappa_{L}>0$ such that $|\tilde{t}_{k}-t_{k}|\leq\tau_{L}c(t_{k})^{\kappa_{L}}$. Given that $c(t_{k})=\exp(1-1/\rho(t_{k}))$, a sufficiently large constant $K_{0}$ ensures $\tau_{L}c(t_{k})^{\kappa_{L}}<t_{k}$ when $k>K_{0}$. It follows that $\tilde{t}_{k}>0$ for all $k>K_{0}$. Consequently, we deduce that $\lim_{k\to\infty}\sigma(\gamma^{*}(t_{k}))=\lim_{k\to\infty}\sigma(\tilde{\gamma}(\tilde{t}_{k}))$, indicating a normal-form perfect equilibrium. This completes the proof. ∎
@@ 4.2 A Smooth Path to a Normal-Form Perfect Equilibrium
In this subsection, we establish the existence of a smooth path consisting of solutions to the system (10). This path starts from an arbitrary initial realization plan and ultimately converge into a normal-form perfect equilibrium.
@@
At $t=2$, the system (10) has a unique solution, given by $(\gamma^{*}(2),\lambda^{*}(2),\nu^{*}(2))$, with the components satisfying $\gamma^{*i}(2;\varpi^{i}_{I^{j}_{i}}a)=\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)$, $\lambda^{i}(2;\varpi^{*i}_{I^{j}_{i}}a)=1$ and $\nu^{*i}_{I^{j}_{i}}(2)=1$.
@@
$\begin{array}[]{l}\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(2;\varpi^{i}_{I^{j}_{i}}a)-\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(2;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ 0<\gamma^{i}(2;\varpi^{i}_{I^{j}_{i}}a),\;0<\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (12)
Suppose that $(\gamma^{*}(2),\lambda^{*}(2),\nu^{*}(2))$ is a solution to the system (12). We proceed by substituting the first set of equations into the third set, and summing over $a\in A(I^{j}_{i})$ for each $i\in N,j\in M_{i}$, obtaining $\nu^{*i}_{I^{j}_{i}}(2)=1$. Consequently, the first set of equations yields $\lambda^{*i}(2;\varpi^{i}_{I^{j}_{i}}a)=\nu^{*i}_{I^{j}_{i}}(2)=1$ for $i\in N,j\in M_{i},a\in A(I^{j}_{i})$. Substituting these results back into the third set, it follows that $\gamma^{*i}(2;\varpi^{i}_{I^{j}_{i}}a)=\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)$ for $i\in N,j\in M_{i},a\in A(I^{j}_{i})$. The proof is complete. ∎
As shown in Lemma 6, the system (10) admits a unique solution at $t=2$. We then identify a connected component containing this solution that extends to intersect the $t=0$ level. To facilitate this, it is crucial to introduce Browder’s fixed-point theorem [Browdercontinuityfixedpoints1960].
@@
Let $\Omega_{L}=\cup_{t\in(0,2]}\Omega(t)$. For $(\hat{\gamma},t)\in\Omega_{L}\times(0,2]$, let $\varphi(\hat{\gamma},t)=(\gamma^{i}(t):i\in N)$, where $\gamma^{i}(t)$ represents the unique solution of the strictly convex optimization problem,
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-c(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}))\\ &+c(t)\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\\ &+\theta(t)\sum\limits_{(j,a)\notin D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a))-\frac{1}{2}\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\hat{\gamma}^{i}(\varpi^{i}_{I^{j}_{i}}a))^{2}\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (13)
For $(\hat{\gamma},0)\in\Omega_{L}\times\{0\}$, define $\varphi(\hat{\gamma},0)=(\gamma^{i}(0):i\in N)$, where $\gamma^{i}(0)$ corresponds to the unique solution of the strictly convex optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(0)}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(0;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i})-\frac{1}{2}\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\gamma^{i}(0;\varpi^{i}_{I^{j}_{i}}a)-\hat{\gamma}^{i}(\varpi^{i}_{I^{j}_{i}}a))^{2}\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(0;\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(0;\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (14)
@@
According to Theorem 2.2.2 in [FiaccoIntroductionSensitivityStability1983], the mapping $\varphi(\gamma,t)$ is continuous from $\Omega_{L}\times[0,2]$ into $\Omega_{L}$. Define the set $\mathscr{F}=\{(\gamma(t),t)\in\Omega_{L}\times[0,2]\mid\varphi(\gamma(t),t)=\gamma(t)\}$. Theorem 4 guarantees the existence of a connected component of $\mathscr{F}$ that intersects both $\mathbb{R}^{n_{0}}\times\{2\}$ and $\mathbb{R}^{n_{0}}\times\{0\}$. We designate this connected component by $\mathscr{F}^{c}$, and specifically identify the subset corresponding to $t>0$ as $\widetilde{\mathscr{F}}^{c}$.